# Gravitational-wave memory effects in the Damour-Esposito-Farèse extension of Brans-Dicke theory

### v1

@@ I Introduction
Specifically, the organization of this paper is as follows: In Sec. II, we review aspects of ST theories in the Einstein frame in Bondi-Sachs coordinates. We use the Einstein-frame results to determine appropriate asymptotically flat boundary conditions and the corresponding form of the metric in the Jordan frame. We then compute the field equations of DEF theory in the Jordan frame and derive asymptotically flat solutions. In Sec. III, we discuss the leading and subleading tensor and scalar memory effects in DEF theory. We then describe how the tensor effects can be computed using the conservation equations for the mass and angular momentum aspects. The scalar effects are computed from charge contributions, and the subleading scalar has a new nonlinear contribution in DEF theory that we discuss. We also describe how the PN inspiral memory signals from compact binaries differ between DEF and BD theories. We present our conclusions in Sec. IV. Finally, we discuss BMS symmetries and the transformation of the Bondi metric functions and scalar fields under these symmetries in Appendix A.
@@
Throughout this paper, we use units in which the speed of light is $c=1$ and in which the gravitational constant in the Einstein frame satisfies $\tilde{G}_{E}=1$. We use the conventions for the metric and curvature tensors given in Misner et al. (1973). Greek indices ($\mu,\nu,\alpha,\dots$) represent four-dimensional spacetime indices, and uppercase Latin indices $(A,B,C,\dots)$ represent indices on the 2-sphere. We used the xAct suite Martín-García (2025, 2008) to perform some of the calculations. A Mathematica notebook with these calculations is available in this GitHub repository Tahura (2025).
@@ II Bondi-Sachs framework for DEF theory
In this section, we review the formulation of DEF theory and give the forms of the field equations in Bondi-Sachs coordinates. We will typically refer to the theory as DEF theory, though we may also refer to it as ST theory, as is done in Ma et al. (2023). First, we give a brief overview of DEF theory in the Einstein frame in Sec. II.1; the form of the field equations and the assumptions that we make for the boundary conditions for asymptotic flatness are identical to those in BD theory Tahura et al. (2020); Hou and Zhu (2020a). Next, we transform to the Jordan frame to obtain the corresponding fall-off conditions of the Bondi metric functions and the scalar field; the transformation is of precisely the same form as in BD theory Tahura et al. (2020). Finally, we compute the field equations in the Jordan frame and derive asymptotically flat solutions in Sec. II.3.
@@ II.1 Einstein-frame field equations and fall-off conditions
We use the notation $\tilde{g}_{\mu\nu}$ for the metric in the Einstein frame, $\tilde{R}$ for the Ricci scalar, and ~$\tilde{\nabla}_{\mu}$ for the covariant derivative compatible with $\tilde{g}_{\mu\nu}$. The field $\Phi$ is a massless real scalar, and we set the “bare” gravitational constant [^1] The phrase “bare” gravitational constant refers to the constant that appears in the action. The locally measured value of Newton’s constant can differ from the bare one in modified theories of gravity. $\tilde{G}_{E}$ in the Einstein frame equal to one. Varying the action with respect to the metric and the scalar field gives the field equations of DEF theory:
@@
We now introduce Bondi coordinates $(\tilde{u},\,\tilde{r},\,\tilde{x}^{A})$, which are well adapted for studying outgoing radiation from an isolated source. The variable $\tilde{u}$ is the retarded time, $\tilde{r}$ is a radial coordinate, and $\tilde{x}^{A}$ are coordinates on two-sphere surfaces of constant $\tilde{u}$ and $\tilde{r}$ ($A$ can be either $1$ or $2$). The explicit form of the metric is (see Bondi et al. (1962); Sachs (1962); Mädler and Winicour (2016))
@@
$+\tilde{r}^{2}\tilde{h}_{AB}\left(d\tilde{x}^{A}-\tilde{U}^{A}d\tilde{u}\right)\left(d\tilde{x}^{B}-\tilde{U}^{B}d\tilde{u}\right).$ (3)
The four Bondi metric functions—$\tilde{V}$, $\tilde{\beta}$, $\tilde{U}^{A}$ and $\tilde{h}_{AB}$, which depend on all four Bondi coordinates—contain six degrees of freedom, because the symmetric tensor $\tilde{h}_{AB}$ contains only two degrees of freedom when the determinant condition of Bondi gauge is imposed: namely,
@@
Note that the field equations in Eq. (2) have the same form as the Einstein-Klein-Gordon system in GR with a massless real scalar field, precisely as in BD theory. We thus assume the same fall-off for $\Phi$ as a power series expansion in $1/\tilde{r}$, with the leading-order piece being constant, as in Tahura et al. (2020):
@@
We also impose the same fall-off conditions on metric functions as in Tahura et al. (2020):
@@
$\tilde{V}={}$ $\tilde{r}+\tilde{V}_{0}(\tilde{u},\,\tilde{x}^{C})+\mathcal{O}\left(\frac{1}{\tilde{r}}\right)\,.$ (7d)
@@
Because the fall-off conditions and the evolution equations have precisely the same form as in BD theory, the solutions to the field equations will also have the same form. For example, they give that $\tilde{\beta}_{1}=0$ and $\tilde{U}_{2}^{A}=-(1/2)\eth_{B}\tilde{c}^{AB}$, where ð$\eth_{B}$ is covariant derivative compatible with $q_{AB}$. A more detailed discussion is given in Tahura et al. (2020). As in Tahura et al. (2020), the detailed form of the equations is not essential for the remaining calculations, because we use the Einstein-frame results to determine the fall-off conditions in the Jordan frame. We will solve the field equations in the Jordan frame rather than transforming the full solution in the Einstein frame to the Jordan frame.
@@ II.2 Conformal transformation to the Jordan frame and Jordan-frame fall-off conditions
Given that the differences between DEF and BD theories arise at quadratic order in the scalar field, it is reasonable that the scalar field has the same asymptotic behavior as in BD theory: namely, it goes as a constant plus a function of the coordinates $\tilde{u}$ and $\tilde{x}^{A}$ normalized by $\tilde{r}$. Note, however, that the leading order piece of $\lambda$ has been chosen to be one here, whereas in Tahura et al. (2020) it was allowed to be an arbitrary constant $\lambda_{0}$. Thus, to compare results here with those in Tahura et al. (2020), one should set $\lambda_{0}=1$ in the results of Tahura et al. (2020). [^2] In ST theories, the value of $\lambda_{0}$ is related to the asymptotic value of the gravitational constant, $G$ by $G=(4+2\omega_{0})/[\lambda_{0}(3+2\omega_{0})]$. To set $G=1$ as $r\rightarrow\infty$, then $\lambda_{0}$ would be different from one, in general (see, e.g., the discussion after Eq. (4.6) of Ref. Mirshekari and Will (2013) and also footnote 5 of Ref. Tahura et al. (2020)). Note, however, that the GR limit is $\omega_{0}\rightarrow\infty$, which leads to $G=1$ with $\lambda_{0}=1$. In this paper, we set $\lambda_{0}=1$ because we do not require the asymptotic value of $G$ to be 1 at infinity. Thus, the asymptotic boundary conditions on the fields in the Jordan frame in DEF theory will be the same as those in BD theory, which are described in more detail in Tahura et al. (2020).
We now remind the reader that when performing the conformal transformation in Eq. (8) to obtain the Jordan-frame metric, the Bondi gauge conditions $g_{\tilde{r}\tilde{r}}=g_{\tilde{r}A}=0$ are satisfied, but the determinant condition is not satisfied: namely the determinant $\det[g_{AB}]=\tilde{r}^{4}\lambda^{-2}q(x^{B})$ is no longer independent of $\tilde{u}$ and its $\tilde{r}$ dependence is no longer an overall multiplicative factor of $\tilde{r}^{4}$. To restore the determinant condition, we use the same transformation used in BD theory in Tahura et al. (2020), where we define new coordinates by
@@
Because we will solve the Jordan-frame field equations in the next section, the emphasis here is not on determining the detailed relationship between the Einstein-frame and Jordan-frame metric quantities and scalar fields (the leading-order results are the same as in BD theory; see Tahura et al. (2020)). Rather, we are just determining what are the expected fall off rates in $r$ of the metric functions in the Jordan frame. For example, by enforcing the determinant condition of the Bondi gauge in the coordinates ($u,\,r,\,x^{A}$), the leading order part of the metric component $g_{uu}$ becomes a function of $u$ and $x^{A}$ (this would not occur if we used a gauge that did not enforce this condition; see Hou and Zhu (2020a)). Because scalar-tensor theories admit a scalar “breathing-mode” polarization of GWs, the time dependence of the leading part of $g_{uu}$ is where this radiative degree of freedom enters the metric if the Bondi determinant condition is enforced; the curvature tensors and other properties of the solution are consistent with our notions of asymptotically flat boundary conditions defined in the earlier parts of this section.
@@ II.3.1 Field equations in Bondi gauge
Because the modified Einstein equations (17a) in DEF theory have the same form as in BD theory when the constant $\omega$ of BD theory is replaced by the function $\omega(\lambda)$, the field equations formally have the same form under this replacement. The explicit form of the field equations in the Jordan frame were given in the Appendix of Tahura et al. (2020). Because there are some subtleties related to integrating the “hypersurface equations” that were not discussed there, we describe this in more detail below.
@@
We do not give the evolution equations (the trace-free part of $\mathcal{G}_{AB}=0$) or the supplementary equations ($\mathcal{G}_{uu}=0$ and $\mathcal{G}_{uA}=0$) in full generality here. The expressions are lengthy, and in the solution with asymptotically flat boundary conditions, we will work at an order in $1/r$ in which there are no differences from the results in BD theory given in Tahura et al. (2020).
@@ II.3.2 Asymptotically flat solutions in the Jordan frame
$\omega=\omega_{0}{}+\frac{\omega_{1}{}}{r}+\mathcal{O}(r^{-2})$ (28)
@@
Given that we will need the evolution equation for $h_{AB}$ at leading and next-to-leading order in $1/r$, where we expect no difference from BD theory, it will be simplest to review the solutions of this equation first (although, as discussed in Tahura et al. (2020), for example, the hypersurface equations generally should be solved before the evolution equations are). As in BD theory, the leading-order in $1/r$ part of the evolution equation for $h_{AB}$ gives that $\partial_{u}c_{AB}=N_{AB}$, where $N_{AB}$ is the news tensor. The news (for short) is the free tensor radiative data in this formalism (i.e., the evolution equations do not constrain the evolution of the shear). The next order in $1/r$ of the evolution equation for $h_{AB}$ gives
@@
In the discussion of the hypersurface equations, it was observed that, in general, solving the scalar-field evolution equation requires solving the $\beta$ and $U^{A}$ hypersurface equations before, and solving the hypersurface equation for $V$ simultaneously. However, given the expected fall-offs of the Bondi metric functions and the scalar field in Eq. (16), then as in BD theory, the leading-order part of the scalar-field equation shows that $\partial_{u}\lambda_{1}$ is unconstrained by the evolution equation (making it the free radiative data, which was referred to as the “scalar news” in Tahura et al. (2020)). Thus, $\lambda_{1}$ does not depend on the solutions to the hypersurface equations. The evolution of $\lambda_{2}$ will, so we postpone the discussion of its solution until after we solve the hypersurface equations.

### v2

@@ I.4 Organization and conventions of this paper
The organization of this paper is as follows: In Sec. II, we review aspects of ST theories in the Einstein frame in Bondi-Sachs coordinates. We use the Einstein-frame results to determine appropriate asymptotically flat boundary conditions and the corresponding form of the metric in the Jordan frame. We then compute the field equations of DEF theory in the Jordan frame and derive asymptotically flat solutions. In Sec. III, we discuss the leading and subleading tensor and scalar memory effects in DEF theory. We then describe how the tensor effects can be computed using the conservation equations for the mass and angular momentum aspects. The scalar effects are computed from charge contributions, and the subleading scalar has a new nonlinear contribution in DEF theory that we discuss. We also describe how the PN inspiral memory signals from compact binaries differ between DEF and BD theories and how our results could be used to study the detection prospects of the memory effects, once NR waveforms in DEF theory are available. We present our conclusions in Sec. IV. Finally, we discuss BMS symmetries and the transformation of the Bondi metric functions and scalar fields under these symmetries in Appendix A.
@@
Throughout this paper, we use units in which the speed of light is $c=1$ and in which the gravitational constant in the Einstein frame satisfies $\tilde{G}_{E}=1$. We use the conventions for the metric and curvature tensors given in Misner et al. (1973). Greek indices ($\mu,\nu,\alpha,\dots$) represent four-dimensional spacetime indices, and uppercase Latin indices $(A,B,C,\dots)$ represent indices on the 2-sphere. We used the xAct suite Martín-García (2025, 2008) to perform some of the calculations. A Mathematica notebook with these calculations is available in a GitHub repository Tahura (2025).
@@ II Bondi-Sachs framework for DEF theory
In this section, we review the formulation of DEF theory and give the forms of the field equations in Bondi-Sachs coordinates. We will typically refer to the theory as DEF theory, though we may also refer to it as ST theory, as is done in Ma et al. (2023). First, we give a brief overview of DEF theory in the Einstein frame in Sec. II.1; the form of the field equations and the assumptions that we make for the boundary conditions for asymptotic flatness are identical to those in BD theory (Paper I and Hou and Zhu (2021a)). Next, we transform to the Jordan frame to obtain the corresponding fall-off conditions of the Bondi metric functions and the scalar field; the transformation is of precisely the same form as Paper I in BD theory. Finally, we compute the field equations in the Jordan frame and derive asymptotically flat solutions in Sec. II.3.
As discussed in more detail in Paper I, we follow this approach, because it is simplest to determine asymptotically flat fall-off conditions in the Einstein frame, where the field equations resemble the Einstein-Klein-Gordon system. However, matter fields follow geodesics of the Jordan frame metric, so it is more straightforward to interpret a GW detector’s response in the Jordan frame. Thus, we found it most straightforward to transform the fall-off conditions on the fields to the Jordan frame and directly solve the field equations in the Jordan frame.
@@ II.1 Einstein-frame field equations and fall-off conditions
We use the notation $\tilde{g}_{\mu\nu}$ for the metric in the Einstein frame, $\tilde{R}$ for the Ricci scalar, and $\tilde{\nabla}_{\mu}$ for the covariant derivative compatible with $\tilde{g}_{\mu\nu}$. The field $\Phi$ is a massless real scalar, and we set the “bare” gravitational constant [^2] The phrase “bare” gravitational constant refers to the constant that appears in the action. The locally measured value of Newton’s constant can differ from the bare one in modified theories of gravity. $\tilde{G}_{E}$ in the Einstein frame equal to one. Varying the action with respect to the metric and the scalar field gives the field equations of DEF theory:
@@
We now introduce Bondi coordinates $(\tilde{u},\,\tilde{r},\,\tilde{x}^{A})$, which are well adapted for studying outgoing radiation from an isolated source. The variable $\tilde{u}$ is the retarded time, $\tilde{r}$ is a radial coordinate, and $\tilde{x}^{A}$ are coordinates on two-sphere surfaces of constant $\tilde{u}$ and $\tilde{r}$ (AA can be either 11 or 22). The explicit form of the metric is (see Bondi et al. (1962); Sachs (1962); Mädler and Winicour (2016))
@@
[…]
@@
Note that the field equations in Eq. (2) have the same form as the Einstein-Klein-Gordon system in GR with a massless real scalar field, precisely as in BD theory. We thus assume the same fall-off for $\Phi$ as a power series expansion in $1/\tilde{r}$, with the leading-order piece being constant, as in Paper I:
@@
We also impose the same fall-off conditions on metric functions as in Paper I:
@@
[…]
@@
Because the fall-off conditions and the evolution equations have precisely the same form as in BD theory, the solutions to the field equations will also have the same form. For example, they give that $\tilde{\beta}_{1}=0$ and $\tilde{U}_{2}^{A}=-(1/2)\eth_{B}\tilde{c}^{AB}$, where $\eth_{B}$ is covariant derivative compatible with $q_{AB}$. A more detailed discussion is given in Paper I. As in Paper I, the detailed form of the equations is not essential for the remaining calculations, because we use the Einstein-frame results to determine the fall-off conditions in the Jordan frame. We will solve the field equations in the Jordan frame rather than transforming the full solution in the Einstein frame to the Jordan frame.
@@ II.2 Conformal transformation to the Jordan frame and Jordan-frame fall-off conditions
Given that the differences between DEF and BD theories arise at quadratic order in the scalar field, it is reasonable that the scalar field has the same asymptotic behavior as in BD theory: namely, it goes as a constant plus a function of the coordinates $\tilde{u}$ and $\tilde{x}^{A}$ normalized by $\tilde{r}$. Note, however, that the leading order piece of $\lambda$ has been chosen to be one here, whereas in Paper I it was allowed to be an arbitrary constant $\lambda_{0}$. Thus, to compare results here with those in Paper I, one should set $\lambda_{0}=1$ in the results of Paper I. [^3] In ST theories, the value of $\lambda_{0}$ is related to the asymptotic value of the gravitational constant, $G$ by $G=(4+2\omega_{0})/[\lambda_{0}(3+2\omega_{0})]$. To set $G=1$ as $r\rightarrow\infty$, then $\lambda_{0}$ would be different from one, in general (see, e.g., the discussion after Eq. (4.6) of Ref. Mirshekari and Will (2013) and also footnote 5 of Paper I). Note, however, that the GR limit is $\omega_{0}\rightarrow\infty$, which leads to $G=1$ with $\lambda_{0}=1$. In this paper, we set $\lambda_{0}=1$ because we do not require the asymptotic value of $G$ to be 1 at infinity. Thus, the asymptotic boundary conditions on the fields in the Jordan frame in DEF theory will be the same as those in BD theory, which are described in more detail in Paper I.
We now remind the reader that when performing the conformal transformation in Eq. (8) to obtain the Jordan-frame metric, the Bondi gauge conditions $g_{\tilde{r}\tilde{r}}=g_{\tilde{r}A}=0$ are satisfied, but the determinant condition is not satisfied: namely the determinant $\det[g_{AB}]=\tilde{r}^{4}\lambda^{-2}q(x^{B})$ is no longer independent of $\tilde{u}$ and its $\tilde{r}$ dependence is no longer an overall multiplicative factor of $\tilde{r}^{4}$. To restore the determinant condition, we use the same transformation used in BD theory in Paper I, where we define new coordinates by
@@
Because we will solve the Jordan-frame field equations in the next section, the emphasis here is not on determining the detailed relationship between the Einstein-frame and Jordan-frame metric quantities and scalar fields (the leading-order results are the same as in BD theory; see Paper I). Rather, we are just determining what are the expected fall off rates in $r$ of the metric functions in the Jordan frame. For example, by enforcing the determinant condition of the Bondi gauge in the coordinates ($u,\,r,\,x^{A}$), the leading order part of the metric component $g_{uu}$ becomes a function of $u$ and $x^{A}$ (this would not occur if we used a gauge that did not enforce this condition; see Hou and Zhu (2021a)). Because scalar-tensor theories admit a scalar “breathing-mode” polarization of GWs, the time dependence of the leading part of $g_{uu}$ is where this radiative degree of freedom enters the metric if the Bondi determinant condition is enforced; the curvature tensors and other properties of the solution are consistent with our notions of asymptotically flat boundary conditions defined in the earlier parts of this section.
@@ II.3.1 Field equations in Bondi gauge
Because the modified Einstein equations (17a) in DEF theory have the same form as in BD theory when the constant $\omega$ of BD theory is replaced by the function $\omega(\lambda)$, the field equations formally have the same form under this replacement. The explicit form of the field equations in the Jordan frame were given in the Appendix of Paper I. Because there are some subtleties related to integrating the “hypersurface equations” that were not discussed there, we describe this in more detail below.
@@
We do not give the evolution equations (the trace-free part of $\mathcal{G}_{AB}=0$) or the supplementary equations ($\mathcal{G}_{uu}=0$ and $\mathcal{G}_{uA}=0$) in full generality here. The expressions are lengthy, and in the solution with asymptotically flat boundary conditions, we will work at an order in $1/r$ in which there are no differences from the results in BD theory given in Paper I.
@@ II.3.2 Asymptotically flat solutions in the Jordan frame
$\omega=\omega_{0}{}+\frac{\omega_{1}{}}{r}+\mathcal{O}(r^{-2}),$ (28)
@@
Given that we will need the evolution equation for $h_{AB}$ at leading and next-to-leading order in $1/r$, where we expect no difference from BD theory, it will be simplest to review the solutions of this equation first (although, as discussed in Paper I, for example, the hypersurface equations generally should be solved before the evolution equations are). As in BD theory, the leading-order in $1/r$ part of the evolution equation for $h_{AB}$ gives that $\partial_{u}c_{AB}=N_{AB}$, where $N_{AB}$ is the news tensor. The news (for short) is the free tensor radiative data in this formalism (i.e., the evolution equations do not constrain the evolution of the shear). The next order in $1/r$ of the evolution equation for $h_{AB}$ gives
@@
In the discussion of the hypersurface equations, it was observed that, in general, solving the scalar-field evolution equation requires solving the $\beta$ and $U^{A}$ hypersurface equations before, and solving the hypersurface equation for $V$ simultaneously. However, given the expected fall-offs of the Bondi metric functions and the scalar field in Eq. (16), then as in BD theory, the leading-order part of the scalar-field equation shows that $\partial_{u}\lambda_{1}$ is unconstrained by the evolution equation (making it the free radiative data, which was referred to as the “scalar news” in Paper I). Thus, $\lambda_{1}$ does not depend on the solutions to the hypersurface equations. The evolution of $\lambda_{2}$ will, so we postpone the discussion of its solution until after we solve the hypersurface equations.